OpenStudy (anonymous):

Use the Quadratic Formula to solve the equation. 2x^2 -7x-5=0

pooja195 (pooja195):

Start by listing your abc values

OpenStudy (alexandervonhumboldt2):

ax^2+bx+c=0

OpenStudy (alexandervonhumboldt2):

formula is: \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

a=2 b=7 c=5

OpenStudy (alexandervonhumboldt2):

b and c have wrong sign

OpenStudy (anonymous):

b=-7 c=-5 ?

OpenStudy (mathmale):

good.

OpenStudy (alexandervonhumboldt2):

yes now substitute all this into the formula

OpenStudy (mathmale):

\[y=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (mathmale):

Note that this produces two results (roots).

OpenStudy (anonymous):